# A multi-frequency spaceborne radar perspective of deep convection

### v1

@@ 3.3.2 Global Distributions
Qualitatively, all echo-top proxies generally highlight the same areas, which matches the IWP signal of convective intensity: central Africa, central Argentina and central United States (Fig. 12 a-f). The differences between the KaPR maps and KuPR maps are found in Fig. 12 g-i. The MAML shows the largest difference over central Africa, with KaPR maxima height more than 0.25 standard deviations higher than Ku-band (Fig. 12g). Comparatively, regions of large IWP over the USA and central Argentina show smaller differences in the MAML between KuPR and KaPR. This implies a difference in the structure of convective profiles of KuPR over central Africa, where the MAML is lower in the atmosphere (i.e., closer to the melting level). We hypothesize this is a result of a characteristically different convective environments over central Africa compared to USA or central Argentina, but this is beyond the focus of this paper. However, this serves to highlight how the differences themselves offer hints about the vertical structure of the convection, a topic worthy of future study. As for the differences between the KaPR 20 dBZ and KuPR 30 dBZ echo-tops (Fig. 12h), in general the high latitude oceans have higher KaPR anomalies where deep convection is far less frequent (Fig. 11a) and continental regions show higher KuPR anomalies. Meanwhile the comparison of the 30 dBZ KaPR to the 40 dBZ KuPR shows that the tropical oceanic regions have generally higher Ka-band echo-top anomalies (Fig. 12i). Thus, if one were to use a Ka-band only radar system and the 30 dBZ echo-top, there might be a systematic bias in characterizing stronger convection over the tropical oceans compared to Ku-band systems. In summary, the spatial correspondence between KaPR and KuPR derived convective properties qualitatively match well.
@@ 4 Summary and Conclusions
    • GPM’s Ka-band (KaPR) reflectivity maximized at lower altitudes than the CPR, often 4-5 km from GPM heightStormTop, but does not linearly correspond well with the height of the KuPR maximum in reflectivity ($R^{2}$ of 0.28).
    • Specific KaPR echo-top values produced reasonable linear fits to specific KuPR echo-top values (e.g., 30 dBZ from KaPR to 40 dBZ from KuPR; $R^{2}>0.7$).
@@
    • Maps of echo-tops from KaPR highlight similar regions as KuPR and coincide well with the areas of strong convection delineated by ice water path.
Note that the conclusions from the coincident dataset between GPM and CloudSat are from a relatively small sample size (5000 samples). Unfortunately, several issues with the CPR spacecraft reduce the total size of the coincident dataset, namely the battery failure in 2011 and the loss of most of the reaction wheels to accurately point the radar in 2020. A similar W-band Doppler radar on EarthCARE, launched in May 2024, offers an opportunity to expand such coincidences between GPM and EarthCARE. Furthermore, the EarthCARE Doppler radar information could be leveraged in new ways. Another caveat of this work is that the primary vehicle of analysis were linear relationships (i.e., linear regression) which were used for the sake of simplicity and brevity. Non-linear methods, like random forests or neural networks, have shown good performance in the atmospheric sciences (e.g., Chase et al., 2022a, 2023, and references therein) and might provide more accurate translations from W-band or Ka-band to Ku-band.
Despite these limitations, the statistical similarity of the KaPR to KuPR in this paper supports the assertion that Ka-band only missions (e.g., INCUS) will indeed be able to provide an important characterization of the global distribution of convection including those of the most intense convective storms. This said, this paper primarily focused on the region of deep convective profiles above the melting level which is the likely location of the strongest convective updrafts (e.g., Varble et al., 2014). The use of a Ka-band only system for rainfall mapping will have the added challenge of strong attenuation with liquid phase hydrometeors, but these issues could potentially be mitigated with incorporation of other constraints (e.g., L’Ecuyer and Stephens, 2002) such as coincident radiometric measurements from microwave radiometers or opportunistic passive microwave temperatures from the radar itself (e.g., Battaglia and Panegrossi, 2020).

### v2

@@ 3.3.2 Global Distributions
Qualitatively, all echo-top proxies generally highlight the same regions of high IWP: central Africa, central Argentina and central United States (Fig. 12 a-f). The differences between the KaPR maps and KuPR maps are found in Fig. 12 g-i. The MAML shows the largest difference over central Africa, with KaPR maxima height more than 0.25 standard deviations higher than that of the Ku-band (Fig. 12g). Comparatively, regions of large IWP over the USA and central Argentina show smaller differences in the MAML between KuPR and KaPR. This implies a difference in the structure of convective profiles of KuPR over central Africa, where the MAML is lower in the atmosphere (i.e., closer to the melting level). We hypothesize this is a result of a characteristically different convective environments over central Africa compared to USA or central Argentina, but this is beyond the focus of this paper. However, this serves to highlight how the differences themselves offer hints about the vertical structure of the convection, a topic certainly worthy of future study and an area of interest to INCUS. As for the differences between the KaPR 20 dBZ and KuPR 30 dBZ echo-tops (Fig. 12h), in general the high latitude oceans have greater KaPR anomalies where deep convection is far less frequent (Fig. 11a) and continental regions show higher KuPR anomalies. Meanwhile the comparison of the 30 dBZ KaPR to the 40 dBZ KuPR shows that the tropical oceanic regions have generally higher Ka-band echo-top anomalies (Fig. 12i). Thus, if one were to use a Ka-band only radar system and the 30 dBZ echo-top, there might be a systematic bias in characterizing stronger convection over the tropical oceans compared to Ku-band systems. In summary, the spatial correspondence between KaPR and KuPR echo-tops match well qualitatively.
@@ 4 Summary and Conclusions
    • GPM’s Ka-band (KaPR) reflectivity maximized at lower altitudes than the CPR, often 4-5 km from GPM heightStormTop, but does weakly linearly relate to the height of the KuPR maximum in reflectivity ($R^{2}$ of 0.47).
    • Specific KaPR echo-top values produced reasonable linear fits to specific KuPR echo-top values (e.g., 30 dBZ from KaPR to 40 dBZ from KuPR; $R^{2}of0.59$).
@@
    • Maps of echo-tops from KaPR highlight similar regions as KuPR and coincide well with the areas of large ice water path.
Note that the conclusions from the coincident dataset between GPM and CloudSat are from a relatively small sample size (5000 samples). Unfortunately, several issues with the CPR spacecraft reduce the total size of the coincident dataset, namely the battery failure in 2011 and the loss of most of the reaction wheels to accurately point the radar in 2020. A similar W-band Doppler radar on EarthCARE, launched in May 2024, offers an opportunity to expand such coincidences between GPM and EarthCARE. Furthermore, the EarthCARE Doppler radar information could be leveraged in new ways. Another caveat of this work is that the primary vehicle of analyses were linear relationships (i.e., linear regression) which were used for the sake of simplicity and brevity. Non-linear methods, like random forests or neural networks, have shown good performance in the atmospheric sciences (e.g., Chase et al., 2022, 2023, and references therin) and might provide more accurate translations from W-band or Ka-band to Ku-band. Lastly a complicating factor to the interpretation of the work presented here is Non-Uniform Beam Filling (NUBF). Mroz et al. (2018) showed that hail cores never fill the entire GPM-DPR field of view. Thus, when hail is present, which is often true in deep convective storms, NUBF is present, complicating the interpretation of the reflectivities and the subsequent retrievals.
Despite these limitations, the statistical similarity between the KaPR and KuPR and the potential linear fits between KaPR echo-top height and KuPR echo-top height in this paper supports the assertion that Ka-band only missions (e.g., INCUS) will be able to provide an important characterization of the global distribution of convection. Furthermore, through the comparison of CFADs of the top $1\%$ tallest storms, the KaPR provides valuable insights into even the most intense convective storms. This said, this paper primarily focused on the region of deep convective profiles above the melting level which is the likely location of the strongest convective updrafts (e.g., Varble et al., 2014). The use of a Ka-band only system for rainfall mapping will have the added challenge of strong attenuation with liquid phase hydrometeors, but these issues could potentially be mitigated with incorporation of other constraints (e.g., L’Ecuyer and Stephens, 2002) such as coincident radiometric measurements from microwave radiometers or opportunistic passive microwave temperatures from the radar itself (e.g., Battaglia and Panegrossi, 2020).